Changes

Jump to navigation Jump to search
240 bytes added ,  17:26, April 23, 2010
no edit summary
Line 7: Line 7:     
For example, the probability of a tossed fair coin coming up heads is &frac12; = 0.5 = 50% because there is one favorable outcome (heads) out of 2 possible outcomes (heads and tails). Another common example is standard deck of 52 playing cards: there are 13 cards in each suit: spades, hearts, clubs and diamonds. After the deck is shuffled thoroughly, the probability that the top card on the deck is a heart is  <math>13 / 52 = 1 / 4 = 0.20</math>, or 20%.
 
For example, the probability of a tossed fair coin coming up heads is &frac12; = 0.5 = 50% because there is one favorable outcome (heads) out of 2 possible outcomes (heads and tails). Another common example is standard deck of 52 playing cards: there are 13 cards in each suit: spades, hearts, clubs and diamonds. After the deck is shuffled thoroughly, the probability that the top card on the deck is a heart is  <math>13 / 52 = 1 / 4 = 0.20</math>, or 20%.
 +
 +
==Theory of Evolution==
 +
The famous mathematician Sir Frederick Hoyle once calculated that the chances of life occurring by chance are 1 in 10 to the power of 50, or 1 in 100,000,000,000,000,000,000,000,000,000,000,000,000,000,000,000,000.
    
==See also==
 
==See also==
162

edits

Navigation menu